OpenStudy (mathmath333):

Three runners A,B and C run a race, with runner A finishing 12m ahead of runner B and 18m ahead of runner C, while runner B finishes 8m ahead of runner C. Each runner travels entire distance at a constant speed. What was the length of the race?

OpenStudy (mathmath333):

A. 36m B. 48m C. 60m D. 72m

OpenStudy (anonymous):

@mathmath333 D.

OpenStudy (mathmath333):

@SithsAndGiggles

OpenStudy (anonymous):

If \(x\) is the length of the race, then\[\frac{x-12}{x-18}=\frac{12}{10}\]

OpenStudy (mathmath333):

got it it is \(\large \dfrac{ x−12}{x−18}=\dfrac{x}{x−8}\) thnks